Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 319     Entry time: 23 Aug 2023, 22:56
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, up to 185 pulses at 564 kHz, 3.7  mJ, beam intensity drop, missing last 10-20 pulses in train sometime

Optical laser delivery

  • PP Laser, 800nm, 50 fs, up to 90 pulses @ 282 kHz

Achievements/Observations

  • Alinment for TR-SFX and PAM
  • Data collection for IHR

Issues

  • Beam intensity was not that stable
  • Strange scattering from amorphous sample and powder diffraction were appeared during data collection
ELOG V3.1.4-7c3fd00